two angles are supplementary. if the larger angle is 30 degrees more than twice the smaller angle. find the measure of each angle.
Answer by rapaljer(4667)

Let x = smaller angle (since it is at the end of the first sentence!)
2x + 30 = larger angle

Supplementary means that the sum of the angles is 180, so
x + 2x + 30 = 180
3x + 30 = 180
3x = 150
x=50 = Smaller angle
2x + 30= 2(50) + 30 = 130 = Larger angle.
Check: See that the sum of the angles is 50 + 130 = 180.

R^2 at SCC